# New possibilities in identification of binary choice models with fixed effects

### v5

@@ 3 Main results
Phrasing the problem in terms of support vector machines makes it easy to see that when all the covariates are discrete, we cannot in general expect to point identify $\beta$ up to scale. To see this, notice that by Pakes and Porter, (2024), the sharp identified set for $\beta$ is $\arg\max_{b}E(Y_{1}-Y_{0}){\rm sgn}(W^{\prime}b)$, which can be written as
@@
Therefore, if ${\rm sgn}(W^{\prime}\tilde{\beta})={\rm sgn}(W^{\prime}\beta)$ with probability one, then $\tilde{\beta}$ is in the identified set. When the covariates are discrete, we can in general move $\beta$ a bit and retain the same ${\rm sgn}(W^{\prime}\beta)$. As illustrated in Figure 2, the black and gray lines both perfectly classify the red and blue dots and thus represent two points in the identified set that are not scalar multiple of each other.
Figure 2: Discrete covariates as a support vector machine
@@ 3.1 The sign saturation condition is sufficient for identification
Following Chamberlain, (2010), we assume that one of the regressors is binary in that it is equal to zero at $t=0$ and is equal to one at $t=1$; in other words, one component of $W=X_{1}-X_{0}$ is one. Thus, without loss of generality, we can partition $W=(Z^{\prime},1)^{\prime}\in\mathbb{R}^{K}$. If all the regressors are continuous, then the result can still be applied because the coefficient for the binary regressor is allowed to be zero. The first main result of this paper is to show that the sign saturation condition (together with additional weak assumptions) is enough to guarantee identification up to scaling. To state the formal result, we first recall the definition of the support of a random variable (or its corresponding probability measure): the support of probability measure $\lambda$ is the smallest closed set $A$ such that $\lambda(A)=1$, e.g., page 227 of Dudley, (2002).
@@
3.2 Extension to multiple discrete covariates
In this subsection, we consider the case of multiple discrete covariates. [^5] I thank Xavier D’Haultfœuille for a discussion of this case. We partition $X_{t}=(X_{(1),t}^{\prime},X_{(2),t}^{\prime})^{\prime}$, where $X_{(1),t}\in\mathbb{R}^{K_{1}}$ are discrete covariates and $X_{(2),t}\in\mathbb{R}^{K_{2}}$ are continuous covariates. We can partition the corresponding $W=X_{1}-X_{0}=(D^{\prime},Z^{\prime})^{\prime}$ with $D=X_{(1),1}-X_{(1),0}$ and $Z=X_{(2),1}-X_{(2),0}$. Let $\mathcal{D}$ denote the support of $D$. We show that identification still holds under a conditional version of sign saturation.
@@ 3.2 Extension to multiple discrete covariates
We can easily see that the determinant of this upper triangular matrix is equal to one. Therefore, the above matrix has full rank, which means that $\mathcal{D}$ contains $K_{1}$ linearly independent elements. Then $\beta$ is identified up to scale if for any $j\in\{1,...,K_{1}\}$, $P(E(Y_{1}-Y_{0}\mid Z,R_{0}=K_{1},R_{1}=j)>0)$ and $P(E(Y_{1}-Y_{0}\mid Z,R_{0}=K_{1},R_{1}=j)<0)$ are strictly positive and the support of $Z$ conditional on $(R_{0},R_{1})=(K_{1},j)$ is convex and has non-empty interior.∎
@@ 3.3 Is the sign saturation condition is also necessary?
It turns out that in the many situations, sign saturation characterizes identifiability. To illustrate this point, we consider the setting studied by Chamberlain, (2010) and assume that $u_{t}$ is independent of $(X,\alpha)$ and is from a known distribution. We show that without sign saturation, identification fails at every point unless the distribution of $u_{t}$ is from a special class.
@@ 4 Implications on marginal effects
Here, we explore an important link between $\beta$ and the treatment effects through the sign of components of $\beta$. Let us consider the setting of Section 3.3: (1) the errors $u_{t}$’s are i.i.d across $t$ with distribution $F(\cdot)$ and are independent of $(X,\alpha)$ and (2) $X_{2,t}$ is binary with $X_{2,t}=\mathbf{1}\{t=1\}$. We can interpret $X_{2,t}$ as a binary treatment; in period $t=0$, no one is treated and in period $t=1$, everyone is treated. For $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}$, the sign of $\beta_{2}$ corresponds to the sign of common measures of treatment effects. For example, the average treatment effect effect at $X_{1,t}=z$ is
$\Delta_{ATE}(z)=P(z^{\prime}\beta_{1}+\beta_{2}+\alpha_{i}\geq u_{t})-P(z^{\prime}\beta_{1}+\alpha_{i}\geq u_{t}),$
@@
Although the magnitude of average marginal effect is often not point identified (see e.g., Honoré and Tamer, (2006) and Chernozhukov et al., (2013)), there is hope that the sign of the effect is point identified, which means that the identified set contains only positive numbers, only negative numbers or only zero. Under Assumption 1, the support of $u_{t}$ is $\mathbb{R}$, which means that ${\rm sgn}(\Delta_{ATE}(z))={\rm sgn}(\beta_{2})$ for any $z$. Hence, identifying ${\rm sgn}(\Delta_{ATE}(z))$ is equivalent to identifying ${\rm sgn}(\beta_{2})$. In the literature, there are also terms such as marginal effects or ceteris paribus effects, e.g., Hoderlein and White, (2012) and Chernozhukov et al., (2015). Under the assumptions of a linear index, the sign of these other measures of treatment effects is also the sign of $\beta_{2}$.
@@
By Theorem 1, the sign saturation condition guarantees point identification of $\beta$ up to scale and thus point identification of ${\rm sgn}(\beta_{2})$. By Theorem 5, when the sign saturation condition fails, we cannot distinguish between $\Delta_{ATE}(z)=0$ and $\Delta_{ATE}(z)>0$ unless $\dot{G}(\cdot)$ is periodic. Therefore, unless $\dot{G}(\cdot)$ is periodic, the sign saturation condition is sufficient and necessary to guarantee point identification of the sign of the marginal effects of $X_{2,t}$.
5 Checking the sign saturation condition
@@
By Lemma 1, $P(\phi(X)\leq 0)=1$ if and only if $P(W^{\prime}\beta\leq 0)=1$. Hence, we define $\rho(q)=E\mathbf{1}\{W^{\prime}q\geq 0\}(Y_{1}-Y_{0})$ for $q\in\mathbb{R}^{K}$. Here, we notice that we can replace $\mathbb{R}^{K}$ with $[-1,1]^{K}$ or any set that contains an open neighborhood of zero. We now state the condition of $P(\phi(X)\leq 0)=1$ in terms of $\rho(\cdot)$, which can be easily estimated.
@@
(1) $E(Y_{1}-Y_{0}\mid X)\leq 0$ almost surely if and only if $\sup_{q\in\mathbb{R}^{K}}\rho(q)\leq 0$.
(2) $E(Y_{1}-Y_{0}\mid X)\geq 0$ almost surely if and only if $\inf_{q\in\mathbb{R}^{K}}\rho(q)\geq 0$.
By Lemma 2, testing $E(Y_{1}-Y_{0}\mid X)\leq 0$ almost surely can be done by testing $\sup_{q\in\mathbb{R}^{K}}\rho(q)\leq 0$. This test amounts to constructing a one-sided confidence interval for $\sup_{q\in\mathbb{R}^{K}}\rho(q)$. A similar strategy can be done for testing $E(Y_{1}-Y_{0}\mid X)\geq 0$ almost surely.
Since the sign saturation condition is split into two hypotheses here, one could consider a Bonferroni-type correction. However, Bonferroni-type procedures can have low power. We now construct one statistic that directly tests the sign saturation condition.
By Lemma 2, $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)=0$, i.e., $P\left(E(Y_{1}-Y_{0}\mid X)\leq 0\right)=1$, is equivalent to $\tau_{1}\leq 0$, where $\tau_{1}=\sup_{q\in\mathbb{R}^{K}}\rho(q)$. Similarly, $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)=0$, i.e., $P\left(E(Y_{1}-Y_{0}\mid X)\geq 0\right)=1$, is equivalent to $\tau_{2}\geq 0$, where $\tau_{2}=\inf_{q\in\mathbb{R}^{K}}\rho(q)$. Therefore, the testing problem of
$H_{0}:\ P\left(E(Y_{1}-Y_{0}\mid X)>0\right)=0\ \text{or}\ P\left(E(Y_{1}-Y_{0}\mid X)<0\right)=0$
versus
$H_{1}:\ P\left(E(Y_{1}-Y_{0}\mid X)>0\right),P\left(E(Y_{1}-Y_{0}\mid X)<0\right)>0$
is equivalent to the testing problem of
$H_{0}:\ \min\{\tau_{1},-\tau_{2}\}\leq 0\quad{\rm versus}\quad H_{1}:\ \min\{\tau_{1},-\tau_{2}\}>0.$ (8)
To construct a test statistic, we consider the natural estimate $\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}$, where $\hat{\tau}_{1}=\sup_{q\in\mathbb{R}^{K}}\hat{\rho}(q)$, $\hat{\tau}_{2}=\inf_{q\in\mathbb{R}^{K}}\hat{\rho}(q)$ and
@@
The next question is how to construct the critical value. It turns out that a one-sided confidence interval for $\min\{\tau_{1},-\tau_{2}\}$ is $[\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-c_{1-\alpha}^{*},1]$, where $c_{1-\alpha}^{*}$ satisfies $P(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)>c_{1-\alpha}^{*})=\alpha$ with $S_{n}(q)=\sqrt{n}(\hat{\rho}_{n}(q)-\rho(q))$. Since $S_{n}(\cdot)$ is a simple empirical process, we can use a nonparametric bootstrap to obtain $c_{1-\alpha}^{*}$. We summarize the test below.
@@
Implement the following procedure for testing $H_{0}$ in (8) with nominal size $\alpha\in(0,1)$
@@
    Compute the test statistic $T_{n}=\sqrt{n}\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}$.
@@
    Repeat the previous step many times and compute $c_{1-\alpha}$, the $(1-\alpha)$ quantile of $\sup_{q\in\mathbb{R}^{K}}\tilde{S}_{n}(q)$.
@@
    Reject $H_{0}$ if $T_{n}>c_{1-\alpha}$.
For multiple time periods, simply replace $\mathbf{1}\{W_{i}^{\prime}q\geq 0\}(Y_{i,1}-Y_{i,0})$ with $\sum_{s>t}\mathbf{1}\{(X_{i,s}-X_{i,t})^{\prime}q\geq 0\}(Y_{i,s}-Y_{i,t})$ in the definition of $\hat{\rho}_{n}$ and similarly in the bootstrap statistic $\tilde{S}_{n}$.
Notice that computing $\hat{\tau}_{1}$ is equivalent to computing the maximum score estimator [^7] Notice that maximizing $\hat{\rho}_{n}(q)$ is the same as maximizing $n^{-1}\sum_{i=1}^{n}{\rm sgn}(W_{i}^{\prime}q)(Y_{i,1}-Y_{i,0})$ because ${\rm sgn}(W_{i}^{\prime}q)=-1+2\cdot\mathbf{1}\{W_{i}^{\prime}q\geq 0\}$. and all the existing computational algorithms and software for the maximum score estimator can be used. Finding $\hat{\tau}_{2}$ also reduces to computing the maximum score estimator once we swap $Y_{1}$ and $Y_{0}$. The above algorithm can be justified by the following result.
@@
Let Assumption 1 hold. Assume that $P(Y_{1}=Y_{0})<1$. Then, under $H_{0}$ in (8),
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}>c_{1-\alpha}\right)\leq\alpha.$
If $H_{0}$ does not hold such that $\min\{\tau_{1},-\tau_{2}\}\geq c_{*}$ for a constant $c_{*}>0$, then
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}>c_{1-\alpha}\right)=1.$
@@
Note that the cubic root asymptotics typically associated with the maximum score estimator (see e.g., Kim and Pollard, (1990) and Cattaneo et al., (2020)) does not arise here. The reason is that $\hat{\tau}_{1}$ is the maximum of $\hat{\rho}_{n}(\cdot)$, rather than the argmax. The cubic root asymptotics of the maximum score estimator is due to the non-standard rate of some terms in the expansion for analyzing the argmax. Fortunately, we do not have to deal with such terms for our purpose.
It is worth pointing out that the asymptotic size is exact in the sense that there exists a data-generating process (DGP) such that it satisfies the null hypothesis (8) and the asymptotic size under this DGP is exactly $\alpha$. For example, one can consider the DGP under which $\rho(q)=0$ for any $q$. If the support of $Z_{i}$ (with $W_{i}=(1,Z_{i}^{\prime})^{\prime}$) is not convex but contains a convex open neighborhood of zero, we can restrict $Z_{i}$ to this neighborhood.
Remark 3.
An alternative test is to construct a one-sided confidence interval for $\tau_{1}$ and then for $\tau_{2}$. However, doing inference simultaneously on both $\tau_{1}$ and $\tau_{2}$ is not straight-forward and might invoke Bonferroni-type methods that sacrifice power. Instead, Algorithm 1 circumvents this problem and gives an asymptotically exact test.
For applied researchers, results in this paper suggest the following steps in practice. We can start with blindly computing the maximum score estimator without worrying about whether the identification fails. Then we check the identification by verifying the sign saturation condition using Algorithm 1; in essence, the test statistic and the critical values are essentially the objective function values of the maximum score estimator and its bootstrapped version. Once we are confident that the model is identified, the estimate we already computed is econometrically justified and the bootstrap procedure in Algorithm 1 can also be used for inference via a test inversion.
In this paper, we do not consider non-standard asymptotics under which $P(E(Y_{1}-Y_{0}\mid X)>0)$ and $P(E(Y_{1}-Y_{0}\mid X)<0)$ are local to zero in some sense. A similar issue arises in pre-tests for the strength of instruments in instrumental variables regressions. Perhaps a simple idea is to employ confidence intervals. For example, the one-sided confidence interval with nominal coverage $1-\alpha$ for $\min\{\tau_{1},-\tau_{2}\}$ is simply $[\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-n^{-1/2}c_{1-\alpha},1]$. Even if we fail to reject $H_{0}$ in (8), we would still like to see a relatively high value of $\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-n^{-1/2}c_{1-\alpha}$ to be comfortable with the identification strength. We will leave a more complete solution to this issue as future research.

### v7

@@ 2 Identification of coefficients
Phrasing the problem in terms of support vector machines makes it easy to see that when all the regressors are discrete, we cannot in general expect to point identify $\beta$ up to scale. To see this, notice that by Pakes and Porter, (2024), the sharp identified set for $\beta$ is $\arg\max_{b}E(Y_{1}-Y_{0}){\rm sgn}(W^{\prime}b)$, which can be written as
@@
Therefore, if ${\rm sgn}(W^{\prime}\tilde{\beta})={\rm sgn}(W^{\prime}\beta)$ with probability one, then $\tilde{\beta}$ is in the identified set. When the regressors are discrete, we can in general move $\beta$ a bit and retain the same ${\rm sgn}(W^{\prime}\beta)$. As illustrated in Figure 2, the black and gray lines both perfectly classify the red and blue dots and thus represent two points in the identified set that are not scalar multiple of each other.
Figure 2: Discrete regressors as a support vector machine
@@ 2.1 The sign saturation condition is sufficient for identification
Following Chamberlain, (2010), we assume that one of the regressors is binary in that it is equal to zero at $t=0$ and is equal to one at $t=1$; in other words, one component of $W=X_{1}-X_{0}$ is one. Thus, without loss of generality, we can partition $W=(Z^{\prime},1)^{\prime}\in\mathbb{R}^{K}$. If all the regressors are continuous, then the result can still be applied because the coefficient for the binary regressor is allowed to be zero, see Corollary 1. The first main result of this paper is to show that the sign saturation condition (together with additional weak assumptions) is enough to guarantee identification up to scaling. To state the formal result, we first recall the definition of the support of a random variable (or its corresponding probability measure): the support of probability measure $\lambda$ is the smallest closed set $A$ such that $\lambda(A)=1$, e.g., page 227 of Dudley, (2002).
@@
We now discuss an important special case with only continuous covariates.
Corollary 1.
Let Assumption 1 hold. If the interior of the support of $X_{1}-X_{0}$ contains $\boldsymbol{0}_{\dim(X_{t})}$ (zero in $\mathbb{R}^{\dim(X_{t})})$ and $\beta\neq\boldsymbol{0}_{\dim(X_{t})}$, then $\beta$ is identified up to scale.
A comparison between Example 1 and Corollary 1 highlights the role of discrete variables in identification. If all the covariates are continuous and the change in covariates contains zero in the support, then we can generally expect identification (up to scale) of $\beta$, regardless of whether the covariates are bounded. In contrast, if the covariates include discrete variables (such as the time fixed effect in Chamberlain, (2010)), we no longer have such a generic guarantee of identification.
2.2 Extension to multiple discrete regressors
In this subsection, we consider the case of multiple discrete regressors. [^3] I thank Xavier D’Haultfœuille for a discussion of this case. We partition $X_{t}=(X_{(1),t}^{\prime},X_{(2),t}^{\prime})^{\prime}$, where $X_{(1),t}\in\mathbb{R}^{K_{1}}$ are discrete regressors and $X_{(2),t}\in\mathbb{R}^{K_{2}}$ are continuous regressors. We can partition the corresponding $W=X_{1}-X_{0}=(D^{\prime},Z^{\prime})^{\prime}$ with $D=X_{(1),1}-X_{(1),0}$ and $Z=X_{(2),1}-X_{(2),0}$. Let $\mathcal{D}$ denote the support of $D$. We show that identification still holds under a conditional version of sign saturation.
@@ 2.2 Extension to multiple discrete regressors
We can easily see that the determinant of this upper triangular matrix is equal to one. Therefore, the above matrix has full rank, which means that $\mathcal{D}$ contains $K_{1}$ linearly independent elements. Then $\beta$ is identified up to scale if for every $j\in\{1,...,K_{1}\}$, $P(E(Y_{1}-Y_{0}\mid Z,R_{0}=K_{1},R_{1}=j)>0)$ and $P(E(Y_{1}-Y_{0}\mid Z,R_{0}=K_{1},R_{1}=j)<0)$ are strictly positive and the support of $Z$ conditional on $(R_{0},R_{1})=(K_{1},j)$ is convex and has non-empty interior.∎
@@ 2.3 Is the sign saturation condition is also necessary?
It turns out that in the many situations, sign saturation characterizes identifiability. To illustrate this point, we consider the setting studied by Chamberlain, (2010) and assume that $u_{t}$ is independent of $(X,\alpha)$ and is from a known distribution. [^4] To establish necessity, it is enough to consider the case in which $u_{t}$ is independent of $(X,\alpha)$. This is similar to showing minimax lower bounds. If identification cannot be guaranteed even with the additional assumption of independence between $u$ and $(X,\alpha)$, then it is definitely not guaranteed without this assumption. We show that without sign saturation, identification fails at every point unless the distribution of $u_{t}$ is from a special class.
@@
Remark 2.
As we have seen, the sign saturation condition guarantees the identification of $\beta$ up to scale. Does it also guarantee the parametric rate for estimation? The answer is no because Theorem 2 of Chamberlain, (2010) shows that the information bound is always zero outside the logistic case, regardless of the identification status. This highlights a key difference between identification and the parametric rate in estimation. The question of identification depends on the sign saturation condition, whereas the root-$n$ estimability depends on the semiparametric efficiency calculation or the functional projection argument developed by Bonhomme, (2012).
@@ 3 Implications on marginal effects
Here, we explore an important link between $\beta$ and the treatment effects through the sign of components of $\beta$. Let us consider the setting of Section 2.3: (1) the errors $u_{t}$’s are i.i.d across $t$ with distribution $F(\cdot)$ and are independent of $(X,\alpha)$ and (2) $X_{2,t}$ is binary with $X_{2,t}=\mathbf{1}\{t=1\}$. We can interpret $X_{2,t}$ as a binary treatment; in period $t=0$, no one is treated and in period $t=1$, everyone is treated. For $\beta=(\beta_{1}^{\prime},\beta_{2})^{\prime}$, the sign of $\beta_{2}$ corresponds to the sign of common measures of treatment effects. For example, the average partial effect effect at $X_{1,t}=z$ is
$\Delta_{APE}(z)=P(z^{\prime}\beta_{1}+\beta_{2}+\alpha\geq u_{t})-P(z^{\prime}\beta_{1}+\alpha\geq u_{t}),$
@@
Although the magnitude of average partial effect is often not point identified (see e.g., Honoré and Tamer, (2006) and Chernozhukov et al., (2013)), there is hope that the sign of the effect is point identified, which means that the identified set contains only positive numbers, only negative numbers or only zero. Under Assumption 1, the support of $u_{t}$ is $\mathbb{R}$, which means that ${\rm sgn}(\Delta_{APE}(z))={\rm sgn}(\beta_{2})$ for any $z$. Hence, identifying ${\rm sgn}(\Delta_{APE}(z))$ is equivalent to identifying ${\rm sgn}(\beta_{2})$. In the literature, there are also terms such as average treatment effects or ceteris paribus effects, e.g., Hoderlein and White, (2012) and Chernozhukov et al., (2015). Under the assumptions of a linear index, the sign of these other measures of treatment effects is also the sign of $\beta_{2}$.
@@
By Theorem 1, the sign saturation condition guarantees point identification of $\beta$ up to scale and thus point identification of ${\rm sgn}(\beta_{2})$. By Theorem 5, when the sign saturation condition fails, we cannot distinguish between $\Delta_{APE}(z)=0$ and $\Delta_{APE}(z)>0$ unless $\dot{G}(\cdot)$ is periodic. Therefore, unless $\dot{G}(\cdot)$ is periodic, the sign saturation condition is sufficient and necessary to guarantee point identification of the sign of the marginal effects of $X_{2,t}$.
4 Assessing the sign saturation condition
@@
By Lemma 1, $P(\phi(X)\leq 0)=1$ if and only if $P(W^{\prime}\beta\leq 0)=1$. Hence, we define $\rho(q)=E\mathbf{1}\{W^{\prime}q\geq 0\}(Y_{1}-Y_{0})$ for $q\in\mathbb{R}^{K}$. Here, we notice that we can replace $\mathbb{R}^{K}$ with $[-1,1]^{K}$ or any set that contains an open neighborhood of zero. It turns out that the sign saturation condition can be written in terms of $\rho(\cdot)$ once we observe
$\rho(\beta)=E\left(\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X)\right)\quad{\rm and}\quad\rho(-\beta)=E\left(\mathbf{1}\{\phi(X)\leq 0\}\cdot\phi(X)\right).$
We give the formal statement below.
@@
$\sup_{q\in\mathbb{R}^{K}}\rho(q)=E\max\left\{\phi(X),0\right\}\ {\rm and}\ \inf_{q\in\mathbb{R}^{K}}\rho(q)=E\min\left\{\phi(X),0\right\}.$
Define $\tau_{*}=\min\{\tau_{1},-\tau_{2}\}$ with $\tau_{1}=\sup_{q\in\mathbb{R}^{K}}\rho(q)$ and $\tau_{2}=\inf_{q\in\mathbb{R}^{K}}\rho(q)$. By Lemma 2, measuring $\tau_{*}$ can give us some indication of whether (or how “well”) the sign saturation condition holds. In particular, the sign saturation fails if and only if $\tau_{*}=0$. From the data, we can construct a one-sided confidence interval for $\tau_{*}$. The natural estimate for $\tau_{*}$ is $\hat{\tau}_{*}=\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}$, where $\hat{\tau}_{1}=\sup_{q\in\mathbb{R}^{K}}\hat{\rho}_{n}(q)$, $\hat{\tau}_{2}=\inf_{q\in\mathbb{R}^{K}}\hat{\rho}_{n}(q)$ and
@@
In the proof of Theorem 6 below, we show that
$\sqrt{n}\left(\hat{\tau}_{*}-\tau_{*}\right)\leq\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)\right)\cdot\mathbf{1}\{\tau_{1}\leq-\tau_{2}\}+\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))\right)\cdot\mathbf{1}\{\tau_{1}>-\tau_{2}\},$
where $S_{n}(q)=\sqrt{n}(\hat{\rho}_{n}(q)-\rho(q))$. By the standard arguments of empirical processes, $S_{n}$ converges to a mean-zero Gaussian process, which means that $\sup_{q\in\mathbb{R}^{K}}S_{n}(q)$ and $\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))$ have the same distribution. Thus, a simple $(1-\alpha)$-confidence interval for $\tau_{*}$ can be obtained once we approximate the $(1-\alpha)$ quantile of $\sup_{q\in\mathbb{R}^{K}}S_{n}(q)$. This motivates the following bootstrapping algorithm.
@@
For a $(1-\alpha)$-confidence interval for $\tau_{*}$:
@@
    Compute the estimate $\hat{\tau}_{*}=\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}$.
@@
    Repeat the previous step many times and compute $c_{1-\alpha}$, the $(1-\alpha)$ quantile of $\sup_{q\in\mathbb{R}^{K}}\tilde{S}_{n}(q)$. [^6] In practice, optimization over $q\in\mathbb{R}^{K}$ can be done by a grid search over $N$ points on the unit sphere $\{v\in\mathbb{R}^{K}:\ \|v\|_{2}=1\}$. The computation seems reasonably fast for a normal sample size. For example, in the setting of Section 5 with $\beta_{2}=1$, $n=5000$ and grid size $N=2000$, computing $\sup_{q}\tilde{S}_{n}(q)$ with 500 bootstrap samples take less than 17 seconds in Matlab on a 2022 MacBook Pro laptop. Further speedups are possible depending on $n$, $N$ and available memory.
@@
[…]
Notice that computing $\hat{\tau}_{1}$ is equivalent to computing the maximum score estimator [^7] Notice that maximizing $\hat{\rho}_{n}(q)$ is equivalent to maximizing $n^{-1}\sum_{i=1}^{n}{\rm sgn}(W_{i}^{\prime}q)(Y_{i,1}-Y_{i,0})$ because ${\rm sgn}(W_{i}^{\prime}q)=-1+2\cdot\mathbf{1}\{W_{i}^{\prime}q\geq 0\}$. and all the existing computational algorithms and software for the maximum score estimator can be used. Finding $\hat{\tau}_{2}$ also reduces to computing the maximum score estimator once we swap $Y_{1}$ and $Y_{0}$. The above algorithm can be justified by the following result.
@@
Let Assumption 1 hold. Assume that $P(Y_{1}=Y_{0})<1$. Then
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}(\hat{\tau}_{*}-\tau_{*})>c_{1-\alpha}\right)\leq\alpha.$
@@
Note that the cube-root asymptotics typically associated with the maximum score estimator (see e.g., Kim and Pollard, (1990) and Seo and Otsu, (2018)) does not arise here. The reason is that $\hat{\tau}_{1}$ is the maximum of $\hat{\rho}_{n}(\cdot)$, rather than the argmax. The cube-root asymptotics of the maximum score estimator is due to the non-standard rate of some terms in the expansion for analyzing the argmax. Fortunately, we do not have to deal with such terms for our purpose.
It is worth pointing out that Algorithm 1 is asymptotically exact in the sense that there exists a data-generating process under which the asymptotic coverage probability is exactly $\alpha$. For example, if $\rho(q)=0$ for any $q$, then $\sqrt{n}(\hat{\tau}_{*}-\tau_{*})=\sup_{q\in\mathbb{R}^{K}}S_{n}(q)$ and $P\left(\sqrt{n}(\hat{\tau}_{*}-\tau_{*})>c_{1-\alpha}\right)\rightarrow\alpha$.
Our results characterize what drives the identification and introduce $\tau_{*}$ as a measure for the identification strength. This theoretical insight motivates a preliminary diagnostic step in applied work. Reporting the confidence interval of $\tau_{*}$ in Algorithm 1 can offer important insight on the identification strength. For example, if this confidence interval contains zero, then one should be concerned about identification failure.
However, testing $\tau_{*}=0$ might not detect all the problematic cases. In practice, the identification might not be clear-cut and additional caution is warranted if we worry about the “weak” identification scenario, which can complicate econometric analysis due to the generic issue of post-selection inference, see e.g., Leeb and Pötscher, (2005).
Fortunately, since $\tau_{*}$ can be learned from the data, one option is to proceed with the identification-based inference (i.e., cube-root asymptotics) only when the identification is strong enough. For instance, one might adopt a robust default inference method and switch to the cube-root approach only when a test fails to reject $H_{0}:\ \tau_{*}\geq\tau_{0}$, where $\tau_{0}>0$ is a pre-specified fixed threshold. This is uniformly valid asymptotically as it rules out $\tau_{*}$ being “local” to zero.
This strategy is conceptually analogous to a common treatment of weak instruments in the linear instrumental variable (IV) models: low correlation between the instrument and the exogenous variable suggests potential identification failure and one solution is to proceed with the classical asymptotics only when this correlation is large enough, e.g., measured by the first-stage $F$-statistic (Stock and Yogo,, 2002).
This raises a natural question: what should this default identification-robust method be? In the IV literature, approaches such as the Anderson-Rubin test have been developed. In our model, one could consider inverting the maximum score criterion function, in a manner similar to the Anderson–Rubin approach. While more refined methods may be possible, they would require substantial additional analysis on asymptotic theory, which lies beyond the current focus on characterizing identification and are left for future research.
5 Numerical illustration
We have seen that $\tau_{*}$ is a measure of the sign saturation condition and thus serves a gauge for identification. Using Monte Carlo simulations, we now illustrate how $\tau_{*}$ relates to the identification strength and the accuracy of the cube-root asymptotics.
Consider the model in (1) with $X_{t}=(\mathbf{1}\{t=0\},X_{2,t})^{\prime}$ for $t\in\{0,1\}$, where $X_{2,t}$ is from the uniform distribution on $[-1,1]$, $u_{t}$ is from the standard normal distribution and $\alpha_{t}=(X_{2,0}+X_{2,1})/2$. Here, $X_{2,0}$, $X_{2,1}$, $u_{0}$ and $u_{1}$ are mutually independent. We set $\beta=(1,\beta_{2})^{\prime}$. Therefore, $W^{\prime}\beta=1+\beta_{2}(X_{2,1}-X_{2,0})$. Clearly, the support of $W^{\prime}\beta$ is $[-2|\beta_{2}|+1,2|\beta_{2}|+1]$, which means that the sign saturation holds if and only if $|\beta_{2}|>0.5$. In Figure 3, we plot $\tau_{*}$ (computed using $\hat{\tau}_{*}$ with $n=10^{8}$) as a function of $\beta_{2}$. It confirms the intuition that $\beta_{2}=0.5$ corresponds to $\tau_{*}=0$, which means that the sign saturation condition fails. Higher value of $\beta_{2}$ corresponds to a higher degree to which the sign saturation condition holds.
Figure 3: Coefficient and identification
[…]
Figure 4: Distribution of $n^{1/3}(\hat{\beta}_{2}-\beta_{2})$ for $\beta_{2}=0.6$ (left) and $\beta_{2}=1$ (right)